A 24.6 ml solution of unknown acid contains 0.0034 mols of acid (5.03 g/L) in solution.
Calculate the molar mass of the solution, and use that to calculate the concentration and
pH of the solution.
1
Expert's answer
2020-05-10T09:22:07-0400

Molarity =

"\frac{moles}{volume(in\ litre)}= \frac{0.0034\times1000}{24.6}\\=0.1278\ M"

Molarity =

"H^+ \ ion \ concentration=0.1278\\pH= log(\frac{1}{H^+})=log(\frac{1}{0.1278})=0.8934"